# Quantitative Ultrasound for Periodontal Soft Tissue Characterization

### v1

@@ 3.2 US data acquisition and parametric imaging
Ultrasound imaging data for both phantom and swine studies were acquired by employing a clinical Mindray ultrasound imaging system (ZS3, Mindray Innovation Center, San Jose, CA, USA) equipped with a high-frequency linear array transducer, now commercially available (L30-8, ZS3, Mindray Inc., San Jose, CA) introduced earlier in Figure 2 (a) and (b). The center frequency of the transducer is 18 MHz, the imaging depth and the lateral field of view commonly used are approximately 15 mm and 13 mm, respectively and the transducer elevational focus is at the depth of 8 mm. The raw RF-data were demodulated, and the envelope detected using a Hilbert transform. B-mode images were reconstructed using the logarithmically compressed envelope data within a dynamic range of 70 dB. The Burr model parameters ($b$ and $l$), and the Nakagami model parameters ($m$ and $\alpha$) were locally estimated in user-defined ROIs within the gingiva and the mucosa for each swine to provide additional information about tissue structure using an estimation kernel approach. The sliding kernel moved across ROIs (along horizontal and vertical directions) with an overlap ratio of 70%. Parameter estimations derived from the kernel at each location within the ROI was assigned to the center of the kernel. Linear interpolations were performed between estimations for adjacent centers to obtain a smoother parametric image. Model parameter estimations were mapped as colored parametric images overlaid on the B-mode images. The optimal kernel (window) size for tissue parametric imaging was obtained from the phantom study.
Figure 2: (a) and (b) High-frequency ultrasound transducer for the intraoral scan, (c) the transducer positioning for mid-facial imaging of a molar tooth within the transverse plane in a swine, (d) zoomed-in view of the transducer with the standoff gel pad, (e) an illustration of oral soft tissues anatomy as a general reference for understanding B-scan structure in (f) for a swine case. Important anatomical structures of periodontal hard and soft tissues are annotated in both (e) and (f).
@@ 3.2.1 Optimal window size
In statistical estimations of the Burr and Nakagami parameters for parametric images, there is a trade-off between the spatial resolution of local parameter estimation and statistical variability in estimated parameters. If the window size (WS) is too small, the speckle within the kernel provides unrobust estimations, i.e., the parameters don’t converge. On the other hand, the size of respective tissues under study imposes limitations on the size of the associated ROI and consequenty, how large the window can be. The gingival and alveolar mucosal tissues are inherently small, often about 2 mm thick, which limits the ROI size for the QUS analysis. This trade-off was investigated by evaluating model parameters as a function of WS. Phantom scan envelope data were used to create parametric images of rectangular ROIs (5 mm by 2.5 mm) located in the center of the field of view. WS ranged from 2 to 18 wavelengths with an interval of 2 wavelengths. All data processing in this study was performed using MATLAB (R2023a, MathWorks Inc., Natick MA, USA).
@@ 3.2.2 ROI selection criteria for outlining gingival and mucosal tissues
For ROI selection, the largest possible ROIs were selected for each tissue, excluding regions of hard tissues such as bone, crown as well as the epithelial layer and rete pegs, as those will affect the QUS parameter estimations representing gingiva and mucosa characteristics.
@@ 4 Results
Figure 4: Nakagami parametric imaging for varying window size (same phantom study as shown in Figure 3). In each pair, the left image shows Nakagami $m$ and the right one shows Nakagami $\alpha$ parameter local estimations. All axes are shown in millimeters.
As a note on the optimal WS selection from a practical standpoint of obtaining parametric image for relatively small-sized periodontal soft tissues, the WS of 10 wavelengths reasonably meets criteria of accuracy and tissue size limitation. Therefore, it is suggested as the optimal kernel size for the Burr and Nakagami parametric imaging of swine tissues.
Figure 5: Errorbar plots showing the effect of WS on the Burr parameters (top row), and on the Nakagami paramaters (bottom row) as obtained from an ROI size of 5 mm by 2.5 mm in Figure 3 and Figure 4.
@@ 4.1 Parametric imaging for in vivo swine scans
The parametric images of the Burr and Nakagami parameters of periodontal soft tissues were obtained in ultrasound scans of 10 swine at the baseline condition (no inflammation) using the optimal sliding WS of 10 wavelengths. The results were statistically compared in gingiva versus alveolar mucosa within reasonable ROIs selected for each scan. An example of parametric image of the Burr $b$, Burr $l$, Nakagami $m$ and Nakagami $\alpha$ overlaid on corresponding B-scans are shown in Figure 6 with the reference B-scan. The ROIs selected for these parametric images include the marginal and attached gingiva, alveolar mucosa, epithelium as well as muscle tissues and is meant to show the variations of estimated parameters over the whole scanned region. However, for the quantitative characterization of gingiva and alveolar mucosa, individual ROIs are selected for each tissue excluding any bone, epithelium layers or other regions not associated with these tissues.
Figure 6: Parametric imaging of the Burr and Nakagami parameters in periodontal soft tissues in a swine scan using the WS of 10 wavelengths, for maxilla left first molar tooth. (a) reference B-scan, (b) Burr power-law parameter $b$ and (c) Burr scale factor $l$, (d) Nakagami shape parameter $m$ and (e) Nakagami scale factor $\alpha$.
An example of the ROI selection from these two tissue types is shown in Figure 7 and Figure 8 for the Burr and Nakagami parametric imaging. In these two figures, top rows represent ROI for the gingiva and bottom rows show alveolar mucosal ROI, with the reference B-scan is presented in (e). It is observed that for the gingival tissue, the estimated Burr power-law parameter $b$ and the Nakagami shape parameter $m$ map to a brighter (yellow) colormap, corresponding to higher estimated values compared to mucosal tissues. On the other hand, the Burr scale factor $l$ and the Nakagami scale factor $\alpha$ for gingival tissues are lower than mucosal tissues based on the colormap comparison.
@@
Figure 8: Parametric imaging of the Nakagami parameters in the gingiva (top row) and the alveolar mucosa (bottom row) in a swine scan using the WS of 10 wavelengths. (a) and (c): Nakagami $m$, (b) and (d): Nakagami $\alpha$. The reference B-scan in shown in (e).
@@ 5 Discussion
5.1 Statistical analysis of swine populations
To characterize these two types of periodontal soft tissues adjacent to each other at the baseline condition (no inflammation) for all swine populations, the summary of QUS analyses are presented as boxplots for the Burr and Nakagami models in Figure 9 and Figure 10, respectively. In these figures, $p-values$ are reported to compare statistical significance between the two tissue types. In each boxplot, the blue line shows statistical median for the population and the pentagon symbols represents the statistical mean of the estimations.
In Figure 9, (a) and (b) shows the Burr $b$ and Burr $l$ results. For the Burr parameters, $p-values<0.0001$, indicating that $b$ and $l$ could show statistically significant populations when comparing gingival tissues with alveolar mucosa. The Burr $b$ for gingiva is reported to be higher than mucosa ($b_{Gingiva}=6.6\>(4.8|10.4)$, $b_{Mucosa}=3.6\>(3.2|4.7)$), while the Burr $l$ is lower in gingival tissues ($l_{Gingiva}=254.0\>(177.1|362.7)$, $l_{Mucosa}=851.8\>(571.6|1174.3))$. It is noted that the median (and almost the mean) of the Burr power-law parameter $b$ for the alveolar mucosa tends to be close to the range for normal soft tissues such as liver reported in the literature [22, 33] while this is elevated for gingival tissues.
In Figure 10 for the Nakagami model, (a) shows that Nakagami $m$ is statistically higher for gingiva population in comparison to the alveolar mucosa ($m_{Gingiva}=1.29\>(1.10|1.52)$, $m_{Mucosa}=0.83\>(0.50|1.04))$ and, (b) represents a statistically elevated Nakagami $\alpha$ in alveolar mucosa compared to the gingiva ($\alpha_{Gingiva}=1.82\times 10^{4}\>(0.81\times 10^{4}|3.57\times 10^{4})$, $\alpha_{Mucosa}=4.39\times 10^{5}\>(2.29\times 10^{5}|8.50\times 10^{5}))$. Nakagami $\alpha$ clearly demonstrates that echo intensity in gingiva is significantly lower (an order of magnitude).
Figure 9: Burr parameters for classification of gingiva vs. alveolar mucosa, with boxplots summarizing the average estimations within ROIs in swine cases. (a) Burr $b$, (b) Burr $l$.
Figure 10: Nakagami parameters for classification of gingiva vs. alveolar mucosa, with boxplots summarizing the average estimations within ROIs in swine cases. (a) Nakagami $m$, (b) Nakagami $\alpha$.
Therefore, these plots imply that the Burr and Nakagami parameters show sensitivities to periodontal soft tissue types of gingivae and mucosa and hold potentials to characterize them.
5.2 Histology insight
To cast some insight into a possible explanation for statistically distinct QUS parameters for alveolar mucosa and gingiva, we consulted their histology images. The results for Masson’s Trichrome and H&E stains are shown in Figure 11 and Figure 12, respectively. In these figures, (a) represents histology insights (4x magnification) of the stained tissue slice incorporating both soft and hard tissues. (b) and (c) show a 20x magnified image of gingival and alveolar mucosal tissues, respectively, marked with red dashed box in (a). Scale bars are also to all images for references.
Figure 11: (a) Histology image using Masson’s Trichrome stain (4x magnification). (b) and (c) are enlarged views comparing gingival and alveolar mucosal regions, respectively (20x magnification). Swine oral site: left mandibular first molar.
Figure 12: (a) Histology image using H&E stain (4x magnification) at the same tissue section as shown in Figure 11. (b) and (c) are enlarged views comparing gingival and alveolar mucosal regions, respectively (20x magnification). Swine oral site: left mandibular first molar.
Looking at histology images from both staining techniques, we notice a denser stain in gingival regions compared to alveolar mucosa, which could suggest denser scattering sites in gingiva. This observation from histology images might align with statistical findings from our QUS analyses in two aspects, as outlined below.
@@ 5.2 Histology insight
    First, lower B-scan echo intensity in gingiva compared to alveolar mucosa (indicated by decreased Burr and Nakagami scale factors) is hypothesized to arise from occurrence of multiple scattering between densely-packed scattering sites in gingiva. This would result in a weaker backscatter signal from this region back to the transducer. Thus, gingiva would appear less echogenic in US B-scans compared to alveolar mucosa.
@@
    Second, the histology finding may suggest the presence of higher densities of small, densely-packed scatterers. This could explain elevated estimations of Burr power-law parameter $b$ (associated with the number density of scatterers) obtained for the gingiva compared to the alveolar mucosa in the QUS analysis. Additionally, the QUS analysis showed a higher Nakagami shape parameter $m$ in gingiva compared to alveolar mucosa, indicating a transition of scattering statistics to Rayleigh and post-Rayleigh regime for gingiva. The Rayleigh scattering regime is characterized by the presence of many random small scatterering sites, resulting in a diffuse scattering. Thus, the elevated Nakagami shape parameter in gingiva could be associated with the denser stain observed in the gingiva on histology images, corresponding to an increase in scattering number density, (transitions into the Rayleigh and post-Rayleigh regime).
This histology findings may support our QUS results.
The PDF of the US speckle data within ROIs from alveolar mucosa and gingiva are fitted to the Rayleigh distribution in Figure 13 as an example of PDF comparison. It is observed that speckle statistics of the gingiva fits relatively better to the Rayleigh ($R-squared=0.95$) compared the alveolar mucosa ($R-squared=0.82$). This suggests a gradual transition towards Rayleigh model for gingiva compared to alveolar mucosa. In these PDFs, the Burr model is added as a reference speckle model and it is noteworthy that it is a more accurate fit to statistics of both tissue types with its distinctive heavy-tail behavior at the higher amplitudes compared to the Rayleigh model.
@@
5.3 QUS-based classifications of alveolar mucosa and gingiva
To further investigate the separations of alveolar mucosal and gingival tissues from the QUS standpoint, averages of the Burr and Nakagami parameters are estimated for each swine from local statistical estimations. The results are illustrated as two separate 2D scatterer plots of $l-b$ and $m-\alpha$ in Figure 14 (a) and (b), respectively. In (b), the Nakgami $\alpha$ axis is represented as log-scale to compress the large dynamic ranges of this parameter in alveolar mucosa and gingiva.
In these figures, orange circles represent the estimations for alveolar mucosa and blue diamond symbols show the estimations for gingiva. In each 2D space, a linear boundary is optimized to classify the 2D parameter space by maximizing the accuracy of tissue type prediction. These lines show a clear separation of mucosal and gingival tissues in 2D space of QUS parameters estimated from the US speckle statistics. For the Burr model, the linear classification line gives an accuracy of 93.51%, a sensitivity (true positive rate) of 97.44%, and a specificity (true negative rate) of 89.47% (gingival tissue is assumed to be the positive class). For the Nakagami model, accuracy, sensitivity and specificity are 90.91%, 97.44%, and 84.21%, respectively. For both model, linear classification lines accurately classify all but one gingival case, leading to obtaining similar estimations for sensitivities, however the Burr model estimates more accurate alveolar mucosal cases, resulting in higher specificity that the Nakagami model. By comparing these statistics, the Burr model represents a relatively better separation of the two tissue types, with a slightly higher accuracy and also specificity.
Figure 14: 2D classifications of gingiva vs. alveolar mucosa in swine cases using (a) the Burr model (Burr $b$ vs. Burr $l$), and (b) the Nakagami model (Nakagami $m$ vs. Nakagami $\alpha$). Black lines show linear boundaries between two classes of tissues. Blue diamond symbols: estimations for gingiva, orange circle symbols: estimations for mucosa.
Further, the classification of the two tissue types is investigated by combining the Burr and Nakagami parameters resulting in four features of $b$, $l$, $m$, and $\alpha$, to assess the multi-parametric classification accuracy. To represent the clustering of the two classes with four parameters in a reduced dimensionality (3D) space, the principal component analysis (PCA) is performed on the parameter (feature) space to map the original parameters into a new set of variables, a.k.a. principal components (PCs). PCs are basically directions in the feature space, composed of a linear combination of the original features, along which the data shows the most variation (higher variance). The maximum number of PCs in this case is four, however, we employ the first three PCs to visualize data in 3D. These PCs are the most significant representations of variations (dispersion) in the data. The mapped data onto the PC space along with the mapped classification boundaries are shown in Figure 15. The variance of data captured by the PC 1 is 55.31%, which is the direction with the highest dispersion in the data. PC 2 and PC 3 explain 25.37% and 12.38% of the variance in the data, respectively. Therefore, the first three PCs used to visualize the data in 3D PC space represent 93.06% of the variations of the data. The data are standardized before being mapped to the PC space to be scaled. Therefore, the axes range in this figure are different than the range of original parameters. The decision boundary using the four parameters gives the classification accuracy of 92.21%, which is between the two classification accuracies when applying each model separately. This accuracy is not significantly improved comparing with the 2D classifications. One potential reason behind this is the fact that multi-parametric classification is done based on maximizing the boundary margins whereas the 2D classification is performed by focusing on the accuracy itself directly. It is also noted that PCs are less interpretable representation of the data, and they mostly serve as an effective mean to represent the data in lower dimensional space while retaining most essential information from the data. For a classification with higher number of features, the PCA makes training of the classification model more effective.

### v2

@@ 3.2 US Data Acquisition and Parametric Imaging
Ultrasound imaging data for both phantom and swine studies were acquired by employing a clinical Mindray ultrasound imaging system (ZS3, Mindray Innovation Center, San Jose, CA, USA) equipped with a high-frequency linear array transducer, now commercially available (L30-8, ZS3, Mindray Inc., San Jose, CA) introduced earlier in Figure 2 (a) and (b). For ultrasound data acquisition, the scanner employs the Zone SonographyTM Technique, proprietary to Mindray Inc., in which a broad-spectrum ultrasound pulse that is transmitted through individual zones to utilize synthetic aperture imaging and potentially cover the field of view in fewer transmit/receive cycles [52]. This is contrary to the traditional systems where line-by-line data acquisition only allows for a single focus per transmit thus limiting the imaging depth of field, whereas zone transmits perform only lateral focusing on receive. The raw RF-data were demodulated, and the envelope detected using a Hilbert transform. B-mode images were reconstructed using the logarithmically compressed envelope data within a dynamic range of 70 dB. The Burr model parameters ($b$ and $l$), and the Nakagami model parameters ($m$ and $\Omega$) were locally estimated in user-defined ROIs within the gingiva and the mucosa for each swine to provide additional information about tissue structure using an estimation kernel approach. The sliding kernel moved across ROIs (along horizontal and vertical directions) with an overlap ratio of 70%. Parameter estimations derived from the kernel at each location within the ROI was assigned to the center of the kernel. Model parameter estimations were mapped as colored parametric images overlaid on the B-mode images. The optimal kernel (window) size for tissue parametric imaging was obtained from the phantom study.
It is noted that parameter estimations were performed in MATLAB R2023 (MathWorks, Inc., Natick, Massachusetts). To solve the nonlinear equation for estimating the Burr power-law b, equation 3 is rearranged by shifting all terms to the left size and setting the equation to zero. Then, the zeros of the resulting equation are sought for numerically using a nonlinear solver and a reasonable random initial guess. For estimating the remaining model parameters, associated equations are used directly, as outlined in the Theory section.
Figure 2: (a) and (b) High-frequency toothbrush-sized ultrasound transducer for the intraoral scan, (c) the transducer positioning for mid-facial imaging of a molar tooth within the transverse plane in a swine, (d) zoomed-in view of the transducer with the standoff gel pad, (e) an illustration of oral soft tissues anatomy as a general reference for understanding B-scan structure in (f) for a swine case. Important anatomical structures of periodontal hard and soft tissues are annotated in both (e) and (f).
@@ 3.2.1 Optimal Window Size
In statistical estimations of the Burr and Nakagami parameters for parametric images, there is a trade-off between the spatial resolution of local parameter estimation and statistical variability in estimated parameters. If the window size (WS) is too small, the speckle within the kernel provides unrobust estimations, i.e., the parameters don’t converge. On the other hand, the size of respective tissues under study imposes limitations on the size of the associated ROI and consequenty, how large the window can be. The gingival and alveolar mucosal tissues are inherently small, often about 2 mm thick, which limits the ROI size for the QUS analysis. This trade-off was investigated by evaluating model parameters as a function of WS. Phantom scan envelope data were used to create parametric images of rectangular ROIs (5 mm lateral by 2.5 mm axial) located in the center of the field of view at four different imaging scans. Overall results of these four scans are presented as a boxplot for each WS. In the phantom study, we tested kernels with various sizes each being an integer multiple of the imaging wavelength (64.2 µm based on 1,540 m/s and 24 MHz receive frequency), as the size unit for reference. For this analysis, we followed a standard procedure in finding the optimal window size [21, 32]. Nine different window sizes from 2 to 18 wavelengths with an interval of 2 were tested. Data processing was performed using MATLAB (R2023a, MathWorks Inc., Natick MA, USA).
@@ 3.2.2 ROI Selection Criteria for Outlining Gingival and Mucosal Tissues
For ROI selection, the largest possible ROIs were selected for each tissue, excluding regions of hard tissues such as bone, crown as well as the epithelial layer and rete pegs, as those will affect the QUS parameter estimations representing gingiva and mucosa characteristics The ROI selection for the QUS analysis in each scan was checked with specialists and periodontists to make sure ROIs are selected properly for each tissue type. Distinction of the two tissue types adjacent to each other during the ROI selection was based on two factors: spatial location of landmarks such as bone and epithelium as well as visual speckle pattern differences of gingiva and alveolar mucosa. It is also noted that a small spatial gap was assumed between ROIs in gingiva and alveolar mucosa during ROI placement to prevent including any possible overlapping region.
@@ 4 Results
4.1 Phantom Study
In parametric images, linear interpolations were performed between local estimations of adjacent centers to obtain smoother presentations. This interpolation did not result in any significant variation in statistical parameter estimations. For a selected ROI comparing linear interpolation vs. no interpolation, p-values were assessed using the non-parametric Wilcoxon Rank Sum Test (a.k.a. Mann-Whitney U test), as the data were not normally distributed. Results showed that for the Burr b and l, p-values were 0.29 and 0.28, respectively and for the Nakagami m and $\Omega$, p-values were estimated as 0.82 and 0.57. These results (p-values $>$ 0.05) indicate that the linear interpolation does not impose any statistical significance in the estimated model parameters.
@@ 4.1 Phantom Study
Figure 4: Nakagami parametric imaging for varying window size (same phantom study as shown in Figure 3). In each pair, the left image shows Nakagami $m$ and the right one shows Nakagami $\Omega$ parameter local estimations. All axes are shown in millimeters.
As a note on the optimal WS selection from a practical standpoint where we deal with relatively small-sized periodontal soft tissues, the WS of 10 wavelengths reasonably meets criteria of tissue size limitation and precision. Therefore, it is suggested as the optimal kernel size for the Burr and Nakagami parametric imaging of swine tissues.
Figure 5: Boxplots showing the effect of WS on the Burr (top row), and the Nakagami (bottom row) parameters. Results in boxplots are from four different imaging planes within an ROI size of 5 mm by 2.5 mm (shown in Figure 3 and Figure 4).
@@ 4.2 Parametric Imaging for In vivo Swine Scans
The parametric images of the Burr and Nakagami parameters of periodontal soft tissues were obtained in ultrasound scans of 10 swine at the baseline condition (no inflammation) using the optimal sliding WS of 10 wavelengths. The results were statistically compared in gingiva versus alveolar mucosa within reasonable ROIs selected for each scan. An example of parametric image of the Burr $b$, Burr $l$, Nakagami $m$ and Nakagami $\Omega$ overlaid on corresponding B-scans are shown in Figure 6 with the reference B-scan. The ROIs selected for these parametric images include the marginal and attached gingiva, alveolar mucosa, epithelium as well as muscle tissues and is meant to show the variations of estimated parameters over the whole scanned region. However, for the quantitative characterization of gingiva and alveolar mucosa, individual ROIs are selected for each tissue excluding any bone, epithelium layers or other regions not associated with these tissues.
Figure 6: Parametric imaging of the Burr and Nakagami parameters in periodontal soft tissues in a swine scan using the WS of 10 wavelengths, for maxilla left first molar tooth. (a) reference B-scan, (b) Burr power-law parameter $b$ and (c) Burr scale factor $l$, (d) Nakagami shape parameter $m$ and (e) Nakagami scale factor $\Omega$.
An example of the ROI selection from these two tissue types is shown in Figure 7 and Figure 8 for the Burr and Nakagami parametric imaging. In these two figures, top rows represent ROI for the gingiva and bottom rows show alveolar mucosal ROI, with the reference B-scan is presented in (e). It is observed that for the gingival tissue, the estimated Burr power-law parameter $b$ and the Nakagami shape parameter $m$ map to a brighter (yellow) colormap, corresponding to higher estimated values compared to mucosal tissues. On the other hand, the Burr scale factor $l$ and the Nakagami scale factor $\Omega$ for gingival tissues are lower than mucosal tissues based on the colormap comparison.
@@
Figure 8: Parametric imaging of the Nakagami parameters in the gingiva (top row) and the alveolar mucosa (bottom row) in a swine scan using the WS of 10 wavelengths. (a) and (c): Nakagami $m$, (b) and (d): Nakagami $\Omega$. The reference B-scan in shown in (e).
4.3 Statistical Analysis of Swine Populations
To characterize these two types of periodontal soft tissues adjacent to each other at the baseline condition (no inflammation) for all swine populations, the summary of QUS analyses are presented as boxplots for the Burr and Nakagami models in Figure 9 and Figure 10, respectively. In these figures, $p-values$ are reported to compare statistical significance between the two tissue types. In each boxplot, the blue line shows the statistical median for the population and the pentagon symbols represent the statistical mean of the estimations.
In Figure 9, (a) and (b) shows the Burr $b$ and Burr $l$ results. For the Burr parameters, $p-values<0.0001$, indicating that $b$ and $l$ could show statistically significant populations when comparing gingival tissues with alveolar mucosa. The Burr $b$ for gingiva is reported to be higher than mucosa when comparing mode (Quartile1 |||| Quartile3) of results in gingiva vs. alveolar mucosa ($b_{Gingiva}=6.6\>(4.8|10.4)$, $b_{Mucosa}=3.6\>(3.2|4.7)$). On the other hand, the Burr $l$ is reported as being lower in gingival tissues than in the muscosa ($l_{Gingiva}=254.0\>(177.1|362.7)$, $l_{Mucosa}=851.8\>(571.6|1174.3))$. In Figure 10 for the Nakagami model, (a) shows that Nakagami $m$ is statistically higher for gingiva population in comparison to the alveolar mucosa ($m_{Gingiva}=1.29\>(1.10|1.52)$, $m_{Mucosa}=0.83\>(0.50|1.04))$. (b) represents a statistically elevated Nakagami $\Omega$ in alveolar mucosa compared to the gingiva ($\Omega_{Gingiva}=1.82\times 10^{4}\>(0.81\times 10^{4}|3.57\times 10^{4})$, $\Omega_{Mucosa}=4.39\times 10^{5}\>(2.29\times 10^{5}|8.50\times 10^{5}))$. Nakagami $\Omega$ clearly demonstrates that echo intensity in gingiva is significantly lower (an order of magnitude).
Figure 9: Burr parameters for classification of gingiva vs. alveolar mucosa, with boxplots summarizing the average estimations within ROIs in swine cases. (a) Burr $b$, (b) Burr $l$. In each boxplot, the blue line shows the statistical median for the population and the pentagon symbols represent the statistical mean of the estimations.
Figure 10: Nakagami parameters for classification of gingiva vs. alveolar mucosa, with boxplots summarizing the average estimations within ROIs in swine cases. (a) Nakagami $m$, (b) Nakagami $\Omega$. In each boxplot, the blue line shows the statistical median for the population and the pentagon symbols represent the statistical mean of the estimations.
4.4 Histology Insight
To cast some insight into a possible explanation for statistically distinct QUS parameters for alveolar mucosa and gingiva, we consulted their histology images. The results for Masson’s Trichrome and H&E stains are shown in Figure 12 and Figure LABEL:H&E, respectively. In these figures, (a) represents histology insights (4x magnification) of the stained tissue slice incorporating both soft and hard tissues. (b) and (c) show a 20x magnified image of gingival and alveolar mucosal tissues, respectively, marked with red dashed box in (a). Scale bars are also added to all images for references.
Figure 11: (a) Histology image using Masson’s Trichrome stain (4x magnification). (b) and (c) are enlarged views comparing gingival and alveolar mucosal regions, respectively (20x magnification). Swine oral site: left mandibular.
Figure 12: (a) Histology image using H&E stain (4x magnification). (b) and (c) are enlarged views comparing gingival and alveolar mucosal regions, respectively (20x magnification). Swine oral site: left mandibular.
Looking at histology images from both staining techniques, we notice a denser stain in gingival regions compared to alveolar mucosa, which could suggest denser scattering sites in gingiva.
4.5 QUS-based classifications of alveolar mucosa and gingiva
To further investigate the separations of alveolar mucosal and gingival tissues from the QUS standpoint, averages of the Burr and Nakagami parameters are estimated for each swine from local statistical estimations. The results are illustrated as two separate 2D scatterer plots of $l-b$ and $m-\Omega$ in Figure 13 (a) and (b), respectively. In (b), the Nakgami $\Omega$ axis is represented as log-scale to compress the large dynamic ranges of this parameter in alveolar mucosa and gingiva. In these figures, orange circles represent the estimations for alveolar mucosa and blue diamond symbols show the estimations for gingiva. In each 2D space, a linear boundary is optimized to classify the 2D parameter space by maximizing the accuracy of tissue type prediction. These lines show a clear separation of mucosal and gingival tissues in 2D space of QUS parameters estimated from the US speckle statistics. For the Burr model, the linear classification line gives an accuracy of 93.51%, a sensitivity (true positive rate) of 97.44%, and a specificity (true negative rate) of 89.47% (gingival tissue is assumed to be the positive class). For the Nakagami model, accuracy, sensitivity and specificity are 90.91%, 97.44%, and 84.21%, respectively.
Figure 13: 2D classifications of gingiva vs. alveolar mucosa in swine cases using (a) the Burr model (Burr $b$ vs. Burr $l$), and (b) the Nakagami model (Nakagami $m$ vs. Nakagami $\Omega$). Black lines show linear boundaries between two classes of tissues. Blue diamond symbols: estimations for gingiva, orange circle symbols: estimations for mucosa.
Further, the classification of the two tissue types is investigated by combining the Burr and Nakagami parameters resulting in four features of $b$, $l$, $m$, and $\Omega$, to assess the multi-parametric classification accuracy. To represent the clustering of the two classes with four parameters in a reduced dimensionality (3D) space, the principal component analysis (PCA) is performed on the parameter (feature) space to map the original parameters into a new set of variables, a.k.a. principal components (PCs). PCs are basically directions in the feature space, composed of a linear combination of the original features, along which the data shows the most variation (higher variance). The maximum number of PCs in this case is four, however, we employ the first three PCs to visualize data in 3D. These PCs are the most significant representations of variations (dispersion) in the data. The mapped data onto the PC space along with the mapped classification boundaries are shown in Figure 14. The variance of data captured by the PC 1 is 55.31%, which is the direction with the highest dispersion in the data. PC 2 and PC 3 explain 25.37% and 12.38% of the variance in the data, respectively. Therefore, the first three PCs used to visualize the data in 3D PC space represent 93.06% of the variance of the data. The data are standardized before being mapped to the PC space to be scaled. Therefore, the axes range in this figure are different than the range of original parameters. The decision boundary using the four parameters gives the classification accuracy of 92.21%, which is close to the two classification accuracies when applying each model separately.
@@ 5 Discussion
Our phantom study suggested that WS=10 wavelengths was a reasonable kernel size for local statistical estimation of the Burr and Nakagami parameters, particularly for small-sized periodontal tissues. It is noted that the higher range of variations consistently observed even for the largest WS may be related to the underlying structure of the phantom and might not represent a WS dependence per se. Our findings from QUS analysis of gingiva and alveolar mucosal tissues suggest that Burr and Nakagami parameters show sensitivities to these periodontal tissue types, as supported by Figure 7, and Figure 8 for parametric images and Figure 9 and Figure 10 for the swine population. Thus, QUS hold potentials to characterize periodontal tissues. It is noted that the range of the Burr $b$, which characterize the power-law nature of the scattering structure (scatterer number density), is reported to be close to 3 for some soft tissues such as liver [22, 34]. In this study, the median for the alveolar mucosal tissue tends to be close to this range while this is elevated for gingival tissues.
In this study, our preliminary investigation into multi-parametric characterization of the two tissue types using combined Burr and Nakagami parameters provided a high classification accuracy but close to the 2D classification accuracies. The multi-parametric analysis accuracy did not significantly improve compared with the 2D classifications. The main potential reason behind this is the fact that multi-parametric classification is done based on maximizing the boundary margins whereas the 2D classification is performed by focusing on maximizing the accuracy itself directly rather than separation margins. Another potential explanation is the possible overlap in information explained by the two models. For both model, linear classification lines accurately classify all but one gingival case, leading to obtaining similar estimations for sensitivities, however the Burr model estimates more accurate alveolar mucosal cases, resulting in higher specificity than the Nakagami model. By comparing these statistics, the Burr model represents a relatively better separation of the two tissue types, with a slightly higher accuracy and also specificity. It is also noted that PCs are less interpretable representation of the data, and they mostly serve as an effective mean to represent the data in lower dimensional space while retaining most essential information from the data. For a classification with higher number of features, the PCA makes training of the classification model more effective.
We suggest that QUS findings for periodontal tissue characterization are supported by qualitative analysis of histology images. As introduced in the Introduction, various fiber types exist in the gingiva and alveolar mucosa. They also present in different arrangements with their primary role of protecting the structure of root, bone and soft tissues from degradation and deformation. In ultrasound imaging of periodontal tissues, potential sources of acoustic scattering affecting image formation include fibers as well as vessel walls. Comparing the two tissue types in histology images with statistical QUS findings, observations suggest some consistencies between these findings. Results might align in two aspects, as outlined in following paragraphs.
@@
    First, lower B-scan echo intensity in gingiva compared to alveolar mucosa (indicated by decreased Burr and Nakagami scale factors) is hypothesized to arise from occurrence of multiple scattering between densely-packed scattering sites in gingiva. This could result in a weaker backscatter signal from this region back to the transducer. Thus, gingiva would appear less echogenic in US B-scans compared to alveolar mucosa.
@@
    Second, the histology finding may suggest the presence of higher densities of small, densely-packed scatterers. This could explain elevated estimations of Burr power-law parameter $b$ (associated with the number density of scatterers) obtained for the gingiva compared to the alveolar mucosa in the QUS analysis. Additionally, the QUS analysis showed a higher Nakagami shape parameter $m$ in gingiva compared to alveolar mucosa, indicating a transition of scattering statistics to Rayleigh and post-Rayleigh regime for gingiva. The Rayleigh scattering regime is characterized by the presence of many random small scatterering sites, resulting in a diffuse scattering. Thus, the elevated Nakagami shape parameter in gingiva could be associated with the denser stain observed in the gingiva on histology images, corresponding to an increase in scattering number density, (transitions into the Rayleigh and post-Rayleigh regime). The PDF of the US speckle data within ROIs from alveolar mucosa and gingiva are fitted to the Rayleigh distribution in Figure 15 as an example of PDF comparison. It is observed that speckle statistics of the gingiva fits relatively better to the Rayleigh ($R-squared=0.95$) compared the alveolar mucosa ($R-squared=0.82$). This suggests a gradual transition towards Rayleigh model for gingiva compared to alveolar mucosa. In these PDFs, the Burr model is added as a reference speckle model and it is noteworthy that it is a more accurate fit to statistics of both tissue types with its distinctive heavy-tail behavior at the higher amplitudes compared to the Rayleigh model.
@@
Limitations and Future Direction
In this study, the histology assessment at baseline was based on one sample block of tissues using two different staining methods. In future, we plan to include multiple histology cases and perform statistical and quantitative histology analysis. This will allow a quantitative correlation between histology and QUS. We will investigate histology images through analyzing differences in color densities in gingiva and alveolar mucosa to understand if there is a quantitative correlation between histology findings and QUS parameters in a longitudinal inflammation study. Limitations of this study in terms of the experimental design includes number of porcine for the study and the (still) relatively low resolution of the ultrasound imaging system for scanning small-sized oral tissues. Also, performing QUS analysis on clinical ultrasound data from human subjects should be explored. From the QUS analysis standpoint, future studies should also focus on finding alternate statistical estimators based on higher orders or rational moments of echo envelopes for estimating QUS model parameters and comparing different estimators. Also, the potential effect of WS on the separation accuracy of periodontal tissues could be explored in future. Eventually, future studies should aim at employing other QUS analyses for multi-parametric periodontal diseases classification as well as tissue segmentation.